OpenStudy (anonymous):

solve the equation by completing the square. round to the nearest hundredth if necessary?? x^2 - 6x = -8

jimthompson5910 (jim_thompson5910):

take half of the x coefficient -6 to get -6/2 = -3 then square this to get (-3)^2 = 9

jimthompson5910 (jim_thompson5910):

now add the result of 9 to both sides x^2 - 6x = -8 x^2 - 6x + 9 = -8+9 x^2 - 6x + 9 = 1 what's next?

OpenStudy (anonymous):

Above equation can be rewritten as x^2-6x+8=0 (Now u need to break 8 such that the multiples have their sum =-6 & product =8 Find out factors of 8 for that & try various combinations) Here -4*-2=+8 & -4+-2=-6 Hence Above equation can be written as Hence -2 & -4 are the roots of equation so (x=2 & x=4) are the required solution. Since (x-2)(x-4)=0 (Product is zero if either one or both of them is zero) Alternatively u can always use this formula for solving such equations: ax^2+bx+c=0 then roots are x=[-b+Squrt(b^2-4ac)]/2a & [-b-sqrt(b^2-4ac)]/2a This should help